Calculate the weight of sugar that must be fermented to produce a liter of alcohol (denisty = 0.79 g/mL)


1
Expert's answer
2021-12-10T03:26:30-0500

Solution:

Calculate the mass of alcohol (ethanol):

Mass of ethanol = Density × Volume

Mass of ethanol = (0.79 g/mL) × (1L) × (1000 mL / 1L) = 790 g


Calculate the moles of ethanol:

The molar mass of ethanol is 46.07 g/mol

Therefore,

(790 g) × (1 mol / 46.07 g) = 17.148 mol of ethanol


The fermentation process of sugar (sucrose) into alcohol (ethanol):

C12H22O11(sucrose) + H2O(water) → C6H12O6(glucose) + C6H12O6(fructose)

C6H12O6(glucose) → 2C2H5OH(ethanol) + 2CO2(carbon dioxide)


According to stoichiometry:

Moles of sucrose = Moles of glucose = Moles of ethanol/2

Therefore,

Moles of sucrose = (17.148 mol) / 2 = 8.574 mol


Calculate the mass of sucrose:

The molar mass of sucrose is 342.2965 g/mol

Therefore,

(8.574 mol) × (342.2965 g / 1 mol) = 2934.85 g = 2935 g of sucrose

Mass of sucrose = 2935 g


Answer: 2935 grams of sugar must be fermented